# DGP-LVM: Derivative Gaussian process latent variable models

### v2

@@ 4.5 Results
In Figure 7, we show hyperparameter recovery for the full DGP-LVM in the GP simulation scenario. For most of the simulated datasets, the hyperparameters show good recovery as indicated by low RMSE. We do see some extreme RMSE values though, especially for GP length-scale $\rho$. These extreme cases are explained by the significant increase in the number of estimated parameters, when we consider both varying hyperparameters and correlated outputs without increasing the amount of data. In Figure 8, we show RMSEs when not considering varying hyperparameters or correlated outputs, or both. Clearly, RMSEs are much lower in these cases. We do not display hyperparameter recovery in the periodic simulation scenario as the meaning of true data-generating hyperparameters do not exactly correspond to the hyperparameters of the fitted GP models.
@@ 5 Case study
In order to showcase the application of DGP-LVM to real-world scRNA sequencing data, we re-analyse cell-cycle data from Mahdessian et al., (2021). This dataset comprises of single-cell RNA expression profiles along the cell cycle as well as corresponding RNA velocities as estimates of expression profile derivatives obtained as a pre-processing step of cytopath, a method for simulation based cell trajectory inference (Gupta et al.,, 2022). Briefly, we choose this dataset because it covers single-cell transcriptomic profiles of the cell cycle, i.e. a cyclic process going through four phases depicting substantial variation in gene expressions and velocities.
For the purpose of this case study, we use a reduced data set of spliced RNA gene expression data and its corresponding RNA velocity of 20 cells and 12 genes. We sub-sampled the dataset in a stratified fashion so that cells from all four phases are included. We use the experimental time known as ”cell hours” in the context of this specific data as the prior $\tilde{x}$ for our latent pseudotime (input) $x$. Both $\tilde{x}$ and $x$ are real numbers with values ranging between 0 and 1. For our prior measurement SD, we choose $s=0.03$, so that it is proportional to our choices in simulation studies in Section 4.1.
We fit DGP-LVMs with both our derivative SE and Matern 3/2 covariance functions to showcase the most smooth and the most rough of the common Matern class of covariance functions for which a derivative version exists. For this case study, we use the same weakly informative prior distributions as our simulation study for the GP marginal SD and error SD parameters. For $\rho$ we use a $\text{InvGamma}(5,0.5)$ for the SE and $\text{InvGamma}(5,14)$ for the Matern 3/2 as our GP length scale prior. We these $\rho$ priors, we attempt to account for the varying functional smoothness due to the choice of the covariance function.
@@
Figure 9: Difference of latent pseudotime estimates obtained via DGP-LVM with (a) SE and (b) Matern 3/2 covariance functions and prior cell hours. The point ranges horizontally show 95% prior CIs and vertically show 95% posterior CIs. Notice that the posterior CIs are actually much smaller than the prior CIs since y-axis scale is significantly smaller than x-axis.
As in any real-world latent variable estimation problem, we lack the ground truth to compare the estimated latent values against. Therefore, we study the deviation of the posterior estimates of pseudotime from the cell hours (our prior) by considering the difference or shift in values of the estimated pseudotime from the observed cell hours. The results are shown in Figure 9 with cell hours (prior) on the x-axis and shift (difference of pseudotime and cell hours) on the y-axis. Deviations from the $y=0$ line indicate that latent pseudotimes are different from their cell hours (prior) as a result of learning from gene expression data and velocities. For some cells, prior-posterior differences are up to 5% of the total time scale. Further, we see that the posterior uncertainties (error bars in y-direction) are substantially smaller (considering the scale of the y-axis compared ot x-axis) than the corresponding prior uncertainties (error bars in x-direction), which also indicates that model learning has taken place.
In Figure 10, we show the posterior distributions of GP hyperparameters. We see strongly varying length-scales $\rho$, marginal SDs $\alpha$ and error SDs $\sigma$ across different genes (output dimensions) for both SE and Matern 3/2 models. This clearly points to the necessity of modeling hyperparameters as varying across genes. We also see significant scale differences between the marginal SDs $\alpha$ and $\alpha^{\prime}$ corresponding to gene expression and velocity outputs, respectively. Similar results are seen for the error SDs $\sigma$ and $\sigma^{\prime}$. These results indicate significant scale differences between output RNA gene expression and its derivative RNA velocity. Interestingly, the scale differences go in both directions, such that for some genes $\alpha>\alpha^{\prime}$ and for others $\alpha<\alpha^{\prime}$. While the direction is not important for the DGP-LVM models, it may be highly relevant for understanding the biological processes in which the specific genes are involved.
That said, this case study is meant only as a simple example for demonstrating the application of DGP-LVMs on real-world data. We would like to caution against any specific biological interpretation of the results at this point.
@@
Figure 10: Hyperparameters for DGP-LVM with (a) SE and (b) Matern 3/2 covariance functions. The points indicate posterior mean and the point ranges indicate 95% CIs for each hyperparameter per output dimension. The different colours denote correspondence to the output $y$ or its derivative $y^{\prime}$.
@@ 6 Discussion
Motivated by a real-world problem in the area of single-cell biology, we developed a general statistical modeling framework based on derivative Gaussian processes. In the real-world case, we aim at estimating the latent ordering of cells from RNA gene expression levels and its corresponding time derivative RNA velocity. For this purpose, we developed DGP-LVM, a derivative GP latent variable model that includes derivative information in a multi-output probabilistic framework while accounting for significant scale differences between the outputs and their derivatives. Moreover, it allows the parameters of the model to learn varying information from multiple outputs, which is important in case of scRNA sequencing data where the outputs are different genes, thus accounting for gene-specific information. The model also accounts for between-output correlations since the genes are known to interact with each other. In our extensive simulation studies, we provide clear evidence for the importance of our proposed covariance function modifications. While we specifically focused on modifying the SE and Matern class of covariance functions, our framework is generally applicable for any choice of covariance function that is twice differentiable.
@@
This paper is only the first step towards tackling latent input estimation with derivative Gaussian processes. The current main limitation of DGP-LVMs is their data-scalability as they cannot be easily applied to large amounts of data, such as full sized scRNA sequencing datasets, yet. For a dataset of size $N$, exact GPs have a complexity of $O(N^{3})$ in operations and $O(N^{2})$ in memory. In case of multi-output GPs with correlated outputs and varying hyperparameters, the complexities increase to $O(N^{3}D^{3})$ and $O(N^{2}D^{2})$ respectively, where $D$ is the number of output dimensions (Hensman et al., 2013a, ). Additionally, when performing Bayesian inference via HMC involving a total of $T$ unnormalized log posterior evaluations, the number of operations increases to even $O(N^{3}D^{3}\,T)$. Together, this limits inference for exact GPs on data with large $N$ or $D$. From a real-world data point of view, scRNA sequencing data frequently has a few thousand cells (sample size $N$) with number of genes (output dimensions $D$) in the high hundreds after standard pre-processing steps. In case of DGP-LVM, this issue is even more severe due to adding derivative information, effectively doubling the sample size $N$. To address the computational limitations of DGP-LVM in terms of data-scalability, future research should consider extending approximate GP approaches (e.g., Riutort-Mayol et al.,, 2022) to our DGP-LVM framework.
@@
Another aspect for future research is the choice of prior distributions. Here, we focused on weakly-informative priors for the GP hyperparameters in both our simulation studies and the real-world case study. That said, DGP-LVMs will likely benefit from using stronger priors informed by the application-specific subject matter knowledge, specifically in data-sparse scenarios. This not only applies to priors for the GP hyperparameters, but also to the priors of the latent input variables. The combination of scalable approximations along with stronger prior specifications would foster the general applicability of DGP-LVMs and further increase their accuracy of estimating latent variables.
@@
Figure 11: Summarizing main effects of scaled derivatives for simulated (a) GP data and (b) periodic data for all fitted
Figure 12: Summarizing interaction effects of scaled derivatives with varying hyperparameters for simulated (a) GP data and (b) periodic data
Figure 13: Summarizing interaction effects of scaled derivatives with correlated outputs for simulated (a) GP data and (b) periodic data

### v3

@@ 4.5.2 Model evaluation: Hyperparameters
For most of the simulated datasets, the hyperparameters show good recovery as indicated by low RMSE. We do see some extreme RMSE values though, especially for GP length-scale $\rho$. These extreme cases are explained by the significant increase in the number of estimated parameters, when we consider both varying hyperparameters and correlated outputs without increasing the amount of data.
Figure 8: Matern 3/2 scenario: Hyperparameter RMSEs for (a) full DGP-LVM model and (b) models without scale assumption. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Figure 9: Matern 5/2 scenario: Hyperparameter RMSEs for (a) full DGP-LVM model and (b) models without scale assumption. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Interestingly, we see how the scaling assumption helps identifying the GP marginal SD $\alpha^{\prime}$ and error SD $\sigma^{\prime}$ for the $f^{\prime}$ and $y^{\prime}$ respectively. Without the assumption, the model simply fails to recover the true SD hyperparameters for the derivative part of the data. Additionally, when disabling the assumptions of varying hyperparameters, correlated outputs separately as well as together (see (a), (b) and (c) of Figures 2831 respectively in Appendix B), we see how recovery of hyperparameters, especially the GP marginal SDs struggle due to model misspecification. This demonstrates that each of the model innovations discussed in Section 3 are important when the underlying data require such complexities.
Figure 10: Periodic scenario: Hyperparameter RMSEs for (a) full DGP-LVM model and (b) models without scale assumption. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
4.5.3 Special case: Non-stationary data
In the additional case of periodic data with trend, Figure 11 shows a sharp increase in RMSE values for higher dimensions (when $D=10$). We believe this to be a direct consequence of modelling non-stationary data with stationary GPs. With more number of non-stationary output dimensions, more of the stationary GPs fail to model the data appropriately, thus showing poor model performance in terms of recovering the ground truth of the latent $x$. This special simulation scenario highlights one of the limitations of our current framework, which we further discuss in Section 6.1. Interestingly, for the $D=10$, the problem vanishes when modelling correlated outputs. This is due to the fact that added trend behaviour is same across outputs owing to the shared inputs, thus being highly correlated to one another. Due to this, accounting for correlated outputs seem to improve the recovery of true latent inputs.
Figure 11: Periodic with trend scenario: Main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs
@@ 5 Case study
We showcase the application of DGP-LVM to real-world scRNA sequencing data by re-analysing cell-cycle data from Mahdessian et al., (2021). This dataset comprises of single-cell RNA expression profiles along the cell cycle as well as corresponding RNA velocities as estimates of expression profile derivatives obtained as a pre-processing step of cytopath, a method for simulation based cell trajectory inference (Gupta et al.,, 2022). Briefly, we choose this dataset because it covers single-cell transcriptomic profiles of the cell cycle, i.e. a cyclic process going through four phases depicting substantial variation in gene expressions and velocities.
@@
Figure 12: Difference of latent pseudotime estimates obtained via DGP-LVM with (a) SE and (b) Matern 5/2 covariance functions and prior cell hours. The point ranges horizontally show 95% prior CIs and vertically show 95% posterior CIs. Notice that the posterior CIs are actually much smaller than the prior CIs since y-axis scale is significantly smaller than x-axis.
For the purpose of this case study, we use a reduced data set of spliced RNA gene expression data and its corresponding RNA velocity of 20 cells and 12 genes. In other words, each sample point corresponds to a single cell and each output dimension corresponds to a single gene, with the value being the gene expressions per cell. Thus, for this case study, we have the sample points $N=20$ for $y$ and correspondingly $y^{\prime}$ each with output dimensions $D=12$. We sub-sampled the dataset in a stratified fashion so that cells from all four phases are included. We use the experimental time known as ”cell hours” in the context of this specific data as the prior $\tilde{x}$ for our latent pseudotime (input) $x$. Both $\tilde{x}$ and $x$ are real numbers with values ranging between 0 and 1. For our prior measurement SD, we choose $s=0.03$, so that it is proportional to our choices in simulation studies in Section 4.1.
We fit DGP-LVMs with derivative SE and Matern 5/2 covariance functions. For this case study, We specify the priors for GP marginal SDs $\alpha\sim\text{Normal}^{+}(13.84,3.46^{2})$ and $\alpha^{\prime}\sim\text{Normal}^{+}(1,0.25^{2})$. In case of error SDs we specify $\sigma\sim\text{Normal}^{+}(6.92,3.46^{2})$ and $\sigma^{\prime}\sim\text{Normal}^{+}(0.5,0.25^{2})$. These choices were influenced by the large scaling factor $\lambda$ informed by the data, as the mean and standard deviations of $y$ are on average 13.84 times larger than those of $y^{\prime}$ across the output dimensions. The prior for $\rho$ was specified as $\text{Normal}^{+}(0.4,0.1^{2})$ for the SE and $\text{Normal}^{+}(0.6,0.1^{2})$ for the Matern 5/2 as our GP length scale priors loosely based on the scale of the latent input $x$ (as suggested by the prior $\tilde{x}$). These priors were chosen to account for the varying functional smoothness induced by the choice of covariance function. The DGP-LVM with Matern 3/2 being the least functionally smooth choice of covariance function from the Matern family didn’t converge reasonably with a sensible choice of $\rho$ prior for this specific data and is therefore not presented.
@@
Figure 13: Hyperparameters for DGP-LVM with (a) SE and (b) Matern 5/2 covariance functions. The points indicate posterior mean and the point ranges indicate 95% CIs for each hyperparameter per output dimension. The different colours denote correspondence to the output $y$ or its derivative $y^{\prime}$.
As in any real-world latent variable estimation problem, we lack the ground truth to compare the estimated latent values against. Therefore, we study the deviation of the posterior estimates of pseudotime from the cell hours (our prior) by considering the difference or shift in values of the estimated pseudotime from the observed cell hours. The results are shown in Figure 12 with cell hours (prior) on the x-axis and shift (difference of pseudotime and cell hours) on the y-axis. Deviations from the $y=0$ line indicate that latent pseudotimes are different from their cell hours (prior) as a result of learning from gene expression data and velocities. For some cells, prior-posterior differences are up to 5% of the total time scale. Further, we see that the posterior uncertainties (error bars in y-direction) are substantially smaller (considering the scale of the y-axis compared to x-axis) than the corresponding prior uncertainties (error bars in x-direction), which also indicates that model learning has taken place. Combined with our findings from the simulation study as strong evidence that DGP-LVM is able to learn and recover the true posterior estimates of the latent pseudotimes, these deviations from the prior are interpreted as model learning in the correct direction closer to the true latent ordering of the cells.
In Figure 13, we show the posterior mean along with SD estimates of GP hyperparameters. We see strongly varying length-scales $\rho$, marginal SDs $\alpha$ and error SDs $\sigma$ across different genes (output dimensions) for both SE and Matern 5/2 models. This clearly points to the necessity of modelling hyperparameters as varying across genes. We also see substantial scale differences between the GP marginal SDs $\alpha$ and $\alpha^{\prime}$ corresponding to $f$ and $f^{\prime}$, and consequently gene expression $y$ and velocity $y^{\prime}$ outputs, respectively. Similar, but not as drastic results are seen for the error SDs $\sigma$ and $\sigma^{\prime}$. These results indicate significant scale differences between output RNA gene expression and its derivative RNA velocity. Interestingly, the scale differences go in both directions, such that for some genes $\alpha$ and $\sigma$ are higher than $\alpha^{\prime}$ and $\sigma^{\prime}$. For others, the direction is opposite. While the direction is not important for the DGP-LVM models, it may be highly relevant for understanding the biological processes in which the specific genes are involved. The difference in posterior estimates of the hyperparameters in the different models are heavily influenced by the natural varying functional smoothness corresponding to the choice of different covariance functions.
That said, this case study is meant only as a simple example for demonstrating the application of DGP-LVMs on real-world data. We would like to caution against any specific biological interpretation of the results at this point. The case study was conducted using Apple M1 chip (2 cores in parallel) with 16 GB memory allowances. The runtime (in minutes) for DGP-LVM were approximately 25.3 for the SE model and 73.11 for the Matern 5/2 model.
@@ 6 Discussion
Motivated by a real-world problem in the area of single-cell biology, we developed a class of derivative Gaussian process latent variable models, DGP-LVMs. In the real-world case, we aim at estimating the latent ordering of cells from RNA gene expression levels and its corresponding time derivative RNA velocity. For this purpose, DGP-LVMs not only account for scale differences between the outputs and their derivatives, but also learn from multiple, potentially correlated outputs. The latter is highly important for scRNA sequencing data where the latent cell ordering is informed by many genes, each forming their own output coupled with derivative information.
In our simulation studies, we extensively validate DGP-LVMs demonstrating strong improvements in estimation accuracy of the latent variables by including derivative information. Our results also clearly show the importance of our proposed covariance function modifications. While we specifically focused on modifying the SE and Matern class of covariance functions, our framework is generally applicable for any choice of covariance function that is twice differentiable.
6.1 Limitations and Future Research
@@
This paper is only the first step towards tackling latent variable (input) estimation with derivative Gaussian processes. The current main limitation of DGP-LVMs is their data-scalability as they cannot be easily applied to large amounts of data, such as full sized scRNA sequencing datasets, yet. For a dataset of size $N$, exact GPs have a complexity of $O(N^{3})$ in operations and $O(N^{2})$ in memory. In case of multi-output GPs with correlated outputs and varying hyperparameters, the complexities increase to $O(N^{3}D^{3})$ and $O(N^{2}D^{2})$ respectively, where $D$ is the number of output dimensions (Hensman et al., 2013a, ). Additionally, when performing Bayesian inference via HMC involving a total of $T$ unnormalized log posterior evaluations, the number of operations increases to even $O(N^{3}D^{3}\,T)$. Together, this limits inference for exact GPs on data with large $N$ or $D$. From a real-world data point of view, scRNA sequencing data frequently has a few thousand cells (sample size $N$) with the number of genes (output dimensions $D$) being in the high hundreds after standard pre-processing steps. In case of DGP-LVM, this issue is even more severe due to adding derivative information, effectively doubling the sample size $N$. To address the computational limitations of DGP-LVM in terms of data-scalability, future research should consider extending approximate GP approaches (e.g., Riutort-Mayol et al.,, 2022) to our DGP-LVM framework.
Another limitation of our current DGP-LVMs is their stationary assumption based on the choice of the covariance functions we discuss here. This limits their applicability to non-stationary data as evidenced by our simulation study of periodic data with an added non-linear trend. While this is a general limitation of stationary GPs, the limitation currently lies in not having a derivative version of non-stationary covariance functions. An interesting future research would be to develop DGP-LVMs for non-stationary data where the primary focus would be on obtaining derivative versions of non-stationary covariance functions and verifying their performance for latent variable estimation.
@@
Another aspect for future research is the choice of prior distributions. Here, we focused on informative priors for the GP hyperparameters in both our simulation studies and the real-world case study although they are difficult to come by organically. DGP-LVMs will likely benefit from using stronger priors informed by the application-specific subject matter knowledge, specifically in data-sparse scenarios. This not only applies to priors for the GP hyperparameters, but also to the priors of the latent input variables. Moreover, a joint prior on the covariance function hyperparameters along with latent inputs will likely further improve model convergence. The combination of scalable approximations, improved prior specifications, and additional derivative covariance functions would foster the general applicability of DGP-LVMs, thereby further increasing their ability to accurately estimate latent variables.
@@
Acknowledgments
We acknowledge the Cluster of Excellence iFIT (EXC 2180) ”Image-Guided and Functionally Instructed Tumor Therapies” for supporting Soham Mukherjee. We acknowledge the valuable insights provided by members of ClaassenLab and BürknerLab, Revant Gupta and Marcello Zago. We thank Debapratim Sil and Jayati Chatterjee for their feedback on the manuscript. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Soham Mukherjee.
@@ Acknowledgments
Here we show the mathematical details of the general covariance function structure as well as specific derivative forms of SE, Matern 3/2 and Matern 5/2 covariance functions.
@@
Here we show the additional plots from our simulation studies. Specifically, we provide the MCMC convergence diagnostics for Matern 3/2, 5/2 as well as the periodic simulation scenarios. We then present the full versions of the model evaluation plots for recovery of true latent inputs $x$ using RMSE and MAE as evaluation metrics for all the simulation scenarios. Further we show additional hyperparameter recovery plots based on enabling/disabling different model assumptions. Finally, we present the case where we use four MCMC chains for a reduced SE data simulation scenario that shows minimal or no effect on model evaluation metrics irrespective of number of MCMC chains.
Additional MCMC diagnostic plots
Figure 14: Matern 3/2 scenario: Convergence measures for (a) latent inputs and (b) GP hyperparameters. The individual points correspond to each fitted models per simulated data. The y-axis for Bulk and Tail ESS plots are log10 transformed.
Figure 15: Matern 5/2 scenario: Convergence measures for (a) latent inputs and (b) GP hyperparameters. The individual points correspond to each fitted models per simulated data. The y-axis for Bulk and Tail ESS plots are log10 transformed.
Figure 16: Periodic scenario: Convergence measures for (a) latent inputs and (b) GP hyperparameters. The individual points correspond to each fitted models per simulated data. The y-axis for Bulk and Tail ESS plots are log10 transformed.
Figure 17: Periodic with trend scenario: Convergence measures for (a) latent inputs and (b) GP hyperparameters. The individual points correspond to each fitted models per simulated data. The y-axis for Bulk and Tail ESS plots are log10 transformed.
Full versions of model evaluation plots using RMSE: Latent inputs
Figure 18: Squared exponential scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Figure 19: Matern 3/2 scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Figure 20: Matern 5/2 scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Figure 21: Periodic scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Figure 22: Periodic with trend scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Full versions of model evaluation plots using MAE: Latent inputs
Figure 23: Squared exponential scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.
Figure 24: Matern 3/2 scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.
Figure 25: Matern 5/2 scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.
Figure 26: Periodic scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.
Figure 27: Periodic with trend scenario: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.
Additional model evaluation plots using RMSE:hyperparameters
Figure 28: Squared exponential scenario: Hyperparameter RMSEs for scaled derivatives (a) without varying hyperparameters, (b) without correlated outputs and (c) without both varying hyperparameters and correlated outputs. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Figure 29: Matern 3/2 scenario: Hyperparameter RMSEs for scaled derivatives (a) without varying hyperparameters, (b) without correlated outputs and (c) without both varying hyperparameters and correlated outputs. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Figure 30: Matern 5/2 scenario: Hyperparameter RMSEs for scaled derivatives (a) without varying hyperparameters, (b) without correlated outputs and (c) without both varying hyperparameters and correlated outputs. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Figure 31: Periodic scenario: Hyperparameter RMSEs for scaled derivatives (a) without varying hyperparameters, (b) without correlated outputs and (c) without both varying hyperparameters and correlated outputs. The different colour denotes if the hyperparameters correspond to the original or the derivative part of the model.
Model evaluation plot from simulation study with four MCMC chains
Figure 32: Squared exponential scenario with 4 MCMC chains: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using RMSE.
Figure 33: Squared exponential scenario with 4 MCMC chains: main effects of including (a) scaled derivatives and interaction effects of assuming (b) varying hyperparameters and (c) correlated outputs on recovery of latent inputs (full version) using MAE.